Find the general solution of non-homogenous linear differential equation with constant coefficients.


(D2 - 3D + 2)y = 5e2x

1
Expert's answer
2021-12-09T10:27:11-0500

The auxiliary equation is "m^2-3m+2= 0"

"m^2-m-2m+2=0"

"m(m-1)-2(m-1) = 0"

"(m-1)(m-2) = 0"

"m= 1" or "2"

The solution to the homogeneous part is "y = Ae^x + Be^{2x}"

Suppose the general solution is "y = A(x)e^x + B(x)e^{2x}"

Then, "y'= A'(x)e^x+A(x)e^x+B'(x)e^{2x} + 2B(x)e^{2x}"

Set "A'(x)e^x + B'(x)e^{2x}= 0 ... (*)"

Then "y'" becomes "A(x)e^x+2B(x)e^{2x}"

And "y''= A'(x)e^x+A(x)e^x+2B'(x)e^{2x}+4B(x)e^{2x}"

Substituting all into the given DE we have

"A'(x)e^x+A(x)e^x+2B'(x)e^{2x}+4B(x)e^{2x}-3(A(x)e^x+2B(x)e^{2x})+2(A(x)e^x + B(x)e^{2x}) = 5e^{2x}"

Substituting (*) into the above DE and simplifying further we have

"-A'(x)e^x = 5e^{2x}"

"A'(x)= - 5e^x"

"A(x) =- 5e^x"

And hence "B(x) = 5x"

And the general solution of the DE is

"y = -5e^{2x}+5xe^{2x} = 5(x-1)e^{2x}"